A metal strip is shaped into a right triangle with legs of x and x + 2 and hypotenuse of x + 4. Find the length of the metal strip.
(A) 6 (B) 10 (C) 12 (D) 18 (E) 24

Answer by mananth(16949)

x^2+(x+2)^2=(x+4)^2
x^2 +x^2+4x+4= x^2+8x+16
x^2-4x-12=0
x^2-6x+2x-12=0
x(x-6)+2(x-6)=0
(x-6)(x+2)=0
So x=6
the sides are 6 8 10
length of strip = 24